Assume water boils at 100oC in Houston (near sea level), and at 90oC in Cripple   Creek, Colorado (near 9500 ft ). If it takes 4.8 min to cook an egg in Cripple Creek and 4.5 min in Houston, what is Ea for this process

1
Expert's answer
2021-12-24T04:05:35-0500

Using Arrhenium equation


In "(\frac{k_1}{k_2})=" "-\frac{E_a[\frac{1}{T_1}-\frac{1}{T_2}]}{R}"


"k_1=\frac{1}{4.5}"

"k_2=\frac{1}{4.8}"


"T_1=273+100"

"=373K"


"T_2=273+90"

"=363K"


"R=" Gas constant

"=8.314Joules\/Kelvin .mol"


"In\frac{4.8}{4.5}=\frac{-E_a}{8.314} \begin{bmatrix}
 \frac{1}{373}&-\frac{1}{363}
 
\end{bmatrix}"


"0.064539\u00d78.314=E_a\u00d70.000073856"


"E_a=7265.2Joules\/mol"